Jamie is 4 years younger than her brother. 5 years from now, the sum of their ages will be 32. Find Jamie's present age.
Answer by nerdybill(5404)

Jamie is 4 years younger than her brother. 5 years from now, the sum of their ages will be 32. Find Jamie's present age.
.
Let B = Jamie's Brother's age
then because "Jamie is 4 years younger than her brother" we get
B-4 = Jamie's age
.
From: "5 years from now, the sum of their ages will be 32" we get our formula:
(B+5)+ (B-4+5) = 32
(B+5)+ (B+1) = 32
B+5+B+1 = 32
2B+6 = 32
2B = 26
B = 13 (Jamie's brother age)
.
Jamie's age:
B-4 = 13-4 = 9 years old (Jamie's age)